# A modular perspective to the jet suppression from a small to large radius in very high transverse momentum jets

### v3

@@ Abstract
In this work, we extend the scope of the JETSCAPE framework to cover the jet radius parameter ($R$) dependence of the jet nuclear modification factor, ${R_{AA}}$, for broader area jet cones, going all the way up to $R$ = 1.0. The primary focus of this work has been the in-depth analysis of the high-${p_{T}}$ inclusive jets up to 1 TeV (to probe the quark-gluon plasma medium at much shorter distance scales) and the quenching effects observed in the quark-gluon plasma formed in the Pb-Pb collisions at ${\sqrt{\rm s_{NN}}}$ = 5.02 TeV for the most-central (0-10%) collisions. The nuclear modification factor is calculated for inclusive jets via coupling of the MATTER module (which simulates the high virtuality phase of the parton evolution) with the LBT module (which simulates the low virtuality phase of the parton evolution). These calculations are then compared with the experimental results collected at the ATLAS and the CMS detectors in the jet transverse momentum (${p_{T}}$) ranging from 100 GeV to 1 TeV for the ATLAS detector and greater than 300 GeV up to 1 TeV for the CMS detector. The predictions made by the JETSCAPE are consistent in the high ${p_{T}}$ range as well as for extreme jet cone sizes, with the deviations staying within 10-25%. Our major focus is on calculating the double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=small}}_{\mathrm{AA}}}$) as a function of jet radius and jet-${p_{T}}$, where the observations are well described by the JETSCAPE framework which is based on the hydrodynamic multi-stage evolution of the parton shower.
@@ I Introduction
The extremely hot and dense conditions created at the start of the Big Bang led to what we now understand as the soup of deconfined state of the partons, the quark-gluon plasma (QGP) [1, 2, 3]. The QGP is of such great interest that, to study the properties of this state of matter, experimentalists have spent decades building equipment capable of recreating this extremely dense, soupy state. The Relativistic Heavy-Ion Collider (RHIC) [4, 5] and the Large Hadron Collider (LHC) [6, 7, 8, 9] conduct heavy-ion collisions where the QGP is created for very short instance of time, the parton shower propagation and modification are greatly influenced by the QGP medium. In particular, the high-${p_{T}}$ jets produced in these heavy ion collisions undergo strong yield suppression and medium modification which are together referred to as jet quenching phenomena [10, 11, 12]. Jets modification has been studied in nucleus-nucleus collisions relative to proton-proton collisions to probe the properties of the QGP via constraints from model-to-data comparison  [13, 14, 15, 16]. The measurement of the nuclear modification factor for jets as well as inclusive jet charged-particle fragmentation functions has revealed many important characteristics of the quark-gluon plasma [17, 18, 19, 20]. It provides strong confirmation of the interaction of partons with the deconfined plasma, the respective medium modifications, and the eventual hydrodynamization with the medium [8]. Since there are many conclusive studies based on jet-${R_{AA}}$ [21, 22, 23, 24, 25], our effort here is to push the limits of the current event generators and various energy loss modules for a better description of jet quenching phenomena at very high transverse momenta and broader jet cones with the multi-stage evolution of the parton shower using the Jet Energy-loss Tomography with a Statistically and Computationally Advanced Program Envelope (JETSCAPE) framework (version 3.6.5)  [26].
@@
Along with these measurements, we study inclusive jet spectra for p-p and Pb-Pb by varying resolution parameters in the anti-${k_{T}}$ algorithm which is realized in the FASTJET software package [27]. The inclusive jet spectrum is of significant interest because the sensitivity of hadronization effects is far less than the observables involving individual final-state hadrons. Here, the area of the reconstructed jet cone is defined by the jet radius parameter ${R}$. Thus, by varying ${R}$, the reconstructed jet will include different proportions of energy from the medium response and the quenched jet. We also acknowledge the observations for a new sensitivity to QGP properties and underlying jet quenching mechanism in a study for jet yield suppression versus radius ${R}$ [28]. Specifically, the dependence of jet suppression on jet-${R}$, is predicted by theoretical models based on Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence [29] and perturbative QCD [30].
@@
In this paper, we begin with calculating the jet ${R_{AA}}$ for the Pb-Pb collisions in Section II and compare with the experimental data from the ATLAS as well as the CMS detector for a robust test of configuration and the overall JETSCAPE framework. The calculations include the (2+1)D MUSIC [31] model for hydrodynamics, which is ideal for studying many aspects of heavy-ion collisions. Most notably, this work covers the high-${p_{T}}$ range of jets that is up to 1 TeV, which enables us to probe the QGP medium at much shorter distance scales.
@@
We further exploit the advantages that the JETSCAPE framework offers in Section II.1, that is the coupling of several different energy loss modules such as MARTINI [32] and AdS/CFT [33] with the MATTER [34] to explore the quenching effects in a multi-stage manner. This approach provides insight into the interaction and energy loss mechanisms in the medium concerning the virtuality of the partonic jet. Comparing the experimental data with the trends from these different models, which handle the low virtuality phase, allows us to develop a lucid understanding of the physics governing the above models.
@@
In Section II.2, we proceed with concrete results from the above combinations of several successful models towards the jet radius dependence studies. An important measurement is done by the CMS collaboration with respect to the jet-${R_{AA}}$ in Pb-Pb collisions at ${\sqrt{\rm s_{NN}}}$ = 5.02 TeV at the LHC, for jet cone area covering the radii from 0.2 up to 1  [35]. This measurement gives us an intricate understanding of the behaviour and interaction strength of the high-${p_{T}}$ inclusive collimated jets in proximity to the jet axis, for $R$ = 0.2 and the energy distribution around the cone for values of radii up to 1. The current JETSCAPE framework is based on calculations of perturbative QCD and evolution in a dense medium, thus allowing us to go up to a radius of order 1. We also report the double ratio, i.e., the ratio of ${R_{AA}}$ for a given $R$ with respect to $R$ = 0.2 as a function of jet-${p_{T}}$ and jet radius $R$. This study provides a vivid picture of the energy transactions with the QGP as the jet cone size increases, along with the trends that the JETSCAPE framework predicts.
@@
We conclude this work in Section III, with a concise account of the current JETSCAPE framework’s potential to explain the jet-${R_{AA}}$ for larger area jet cones. We also shed light on the ability of the distinct combination of modules to describe the jet-${R_{AA}}$.
@@
I.1 SIMULATION WITH MULTI-STAGE ENERGY LOSS PERSPECTIVE IN JETSCAPE
@@
The JETSCAPE framework provides an ideal environment for carrying out multi-stage energy loss. The hard scattering is generated by Pythia 8 [36] with initial state radiation (ISR) and multiparton interaction (MPI) [37] turned on, and final state radiation (FSR) turned off. For the event wise simulations, the TRENTo model [38] sets up the initial conditions and the viscous hydrodynamic evolution is described by the (2+1)D MUSIC [31] model. This is then followed by Cooper-Frye prescription [39, 40] which converts the fluid cells to hadrons on an isothermal hypersurface [41] at ${T_{\mathrm{SW}}}$= 151 MeV, where ${T_{\mathrm{SW}}}$ is the temperature of the plasma below which particlization [42] occurs at a certain hypersurface. The jet energy loss induced by scattering is calculated in a succession of two stages: MATTER [34, 43] which takes care of the highly virtual phase (the first stage) while the low virtuality phase (the second stage), is handled concurrently by the LBT model [44, 45, 46]. We have also employed the MARTINI [32, 47, 48] and the AdS/CFT model [33] in combination with the MATTER model to explore the low virtuality phase. The virtuality of the parton is defined as ${Q}^{2}=p^{\mu}p_{\mu}-m^{2}$. The parton undergoes energy loss in two stages, when the virtuality of the parton, ${Q^{2}>Q^{2}_{\mathrm{SW}}}$, where ${Q_{\mathrm{SW}}}$ is the switching virtuality, the MATTER model handles the energy loss and the parton is transferred to the LBT model once ${Q^{2}\leq Q^{2}_{\mathrm{SW}}}$. In these calculations, the jet medium interaction includes inelastic medium-induced gluon radiation and a medium recoil. An extensive account of the comparison of the model to the existing jet-${R_{AA}}$ is already done [49], showing that the contribution of medium recoil is quite significant in the modification of jet-${R_{AA}}$ for a complete set of centrality classes ranging from the most central collisions to the peripheral collisions. This version of the JETSCAPE framework [26] encompasses the modifications of a Hard Thermal Loop (HTL)  [50] for fixed coupling ($\alpha_{\mathrm{s}}^{\mathrm{fix}}$), running coupling ($\alpha_{\mathrm{s}}$), and with a virtuality dependent factor, ${f(Q^{2})}$, that modulates the effective value of jet transport coefficient ($\hat{q}$). The jet transport coefficient $\hat{q}$ is the ratio of the average of the momentum squared (which is exchanged between the jet parton and the medium) to the unit interaction length traversed by the jet parton in a direction perpendicular to the momentum of the jet parton. The JETSCAPE also accounts for the reduced medium-induced emission in the high virtuality phase, due to coherence effects. The calculations based on the HTL considering weak-coupling approximation and the limits of high-temperature yield for a $\hat{q}$ is given as [45],
@@
where $C_{a}$ is the representation specific Casimir, $E$ is the energy of the hard parton, and $T$ is the local temperature of the medium. The coherence effects, which reduce the interaction strength of the parton with the medium are also taken into consideration as the virtuality-dependent modulation factor, which regulates the effective value of $\hat{q}$ in the high virtuality MATTER event generator. The parameterization of the virtuality-dependent modulation factor is given as [49],
@@
here $Q^{2}$ is the running virtuality of the hard parton and $\hat{q}^{\mathrm{run}}_{\mathrm{HTL}}$ denotes the formulation with running coupling. Finally, the partons undergo colorless hadronization according to the default Lund string fragmentation from Pythia 8 [36, 51]. The contributions in final jets are from the hard jet shower part and the effect from the soft medium response, the latter is calculated via the Cooper-Frye formula [39, 40]. We reconstruct jets for several radius selections using the anti-${k_{T}}$ algorithm which is realized in the ${\mathrm{FASTJET}}$ and compare with experimental data. The underlying event  [52, 53] contributions to the jets are minimized by implementing a minimum track requirement of ${p^{\mathrm{track,min}}_{\mathrm{T}}>}$ 0.5 GeV. All the parameters involved in the tuning of the constituent modules follow the standard set of tunes released by the JETSCAPE Collaboration in an elaborate recent study [49].
Figure 1: (Color online) Differential cross-section of inclusive jets for p+p collisions with cone size R = 0.4, with a minimum track requirement of ${p^{\mathrm{track}}_{\mathrm{T}}>}$ 0.5 GeV. Bottom panel shows the JETSCAPE ratio to the ATLAS data [24].
Figure 2: Differential cross-section of inclusive jets in Pb+Pb collisions at ${\sqrt{s_{\mathrm{NN}}}=5.02\rm TeV}$ with cone size $R$ = 0.4, with a minimum track requirement of ${p^{\mathrm{track}}_{\mathrm{T}}>}$ 0.5 GeV. Bottom panel shows the ratio of the JETSCAPE to the ATLAS data [24].
@@ II Results
This paper covers the collision energy: ${\sqrt{s_{\mathrm{NN}}}=5.02}$ TeV for the most central (0-10%) Pb-Pb collisions and includes a comparison with selected experimental data available from the ATLAS and the CMS collaborations for high-${p_{T}}$ jets. The energy loss depicted in this work is achieved by the coupling of MATTER with LBT module. The secondary LBT module (which handles the low virtuality phase) remains the same throughout until and unless specified. Fig. 1 (top panel) shows the ${p+p}$ collision results for inclusive jet spectra at ${\sqrt{s}=5.02~{}\rm TeV}$ for ${|y_{jet}|<2.8}$, which follow the JETSCAPE PP-19 tune [54], are then compared to the experimental data from the ATLAS [24]. In this study, we have set running coupling constant ($\alpha_{\mathrm{s}}$) to 0.35 for all the calculations.
The ratio of inclusive jet cross-section using the JETSCAPE to the ATLAS data is shown in the bottom panel of Fig. 1, which shows that the results are in the acceptable range. We report the 0-10% most central Pb-Pb jet spectra for $R$ = 0.4 in Fig. 2. The ratio of the differential cross-section for inclusive jets in Pb-Pb collisions using JETSCAPE to the ATLAS [24] data are shown in Fig. 2.
@@
Where ${\sigma_{jet}}$ and ${N_{jet}}$ are the inclusive jet cross-section in ${p+p}$ collisions and the jet yield in Pb+Pb collisions, respectively, which are measured as a function of transverse momentum, ${p_{T}}$ and the jet rapidity, $y_{\mathrm{jet}}$. Moreover, ${N_{evt}}$ is the number of Pb+Pb collisions within a specific rapidity interval and $\langle T_{\mathrm{AA}}\rangle$ is the ratio of the mean number of binary nucleon-nucleon collisions $\langle N_{\mathrm{coll}}\rangle$ to the inelastic nucleon-nucleon cross-section ${\sigma_{inel}^{pp}}$, which can be determined from a Glauber model of the nuclear collision geometry . The inclusive jet-${R_{AA}}$ is calculated as the ratio of the Pb-Pb and p-p spectra, which is shown in Fig. 3 in comparison to the ATLAS data [24]. The JETSCAPE (MATTER + LBT) results are in considerable agreement with the actual measurements. However, we see an enhancement (upto ${20\%}$) in the low ${p_{T}}$ region below 150 GeV, which is followed by marginal suppression above 300 GeV, as we ascend in the ${p_{T}}$ range. The effect of the recoil partons in LBT on the total energy loss is not very significant. It is evident that there is a noticeable reduction in both net elastic and radiative jet energy loss when the jet cone includes the recoil partons. These results strongly support the fact that the MUSIC model is not only adequate but also highly effective in describing the experimental observations even at higher jet-${p_{T}}$ in the most central collisions. In contrast to the VISHNU model [55], MUSIC’s hydrodynamic evolution allows for a more accurate representation of the medium’s longitudinal expansion and viscous effects. This leads to a recurring trend of consistent predictions, indicating that MUSIC outperforms other hydrodynamic models in capturing the complex dynamics of jet interactions in the quark-gluon plasma, particularly at high energy scales up to 1 TeV. The MUSIC model shows low sensitivity to coherence effects that dominate at high transverse momentum, compared to VISHNU (2+1)D. This enhances the reliability of MUSIC model for analyzing the jet-suppresions phenomenon at high-${p_{T}}$.
Figure 3: (Color online)the jet-${R_{AA}}$ as a function of jet-${p_{T}}$ for inclusive jets in the most central (0-10%) Pb+Pb collisions at ${\sqrt{s_{\mathrm{NN}}}=5.02\rm~{}TeV}$ for the jet cone radius $R$ = 0.4 compared with the ATLAS data [24].
Figure 4: The jet-${R_{AA}}$ as a function of jet-${p_{T}}$ for inclusive jets in the most central (0-10%) Pb+Pb collisions at ${\sqrt{s_{\mathrm{NN}}}=5.02\rm~{}TeV}$ for the jet cone radius R = 0.2, 0.3, 0.4, 0.6, 0.8 and 1.0 with ${|y_{jet}|<2}$ and a minimum track requirement of ${p^{\mathrm{track}}_{\mathrm{T}}>}$ 0.5 GeV. The plot shows the comparison between the models by red, magenta, and blue markers for (MATTER + LBT), (MATTER + MARTINI), and (MATTER + AdS/CFT), respectively. The predictions are in comparison with CMS data [35] shown with a black marker.

### v4

@@ Abstract
In this work, we expand the scope of the JETSCAPE framework to investigate the dependence of the jet nuclear modification factor, ${R_{AA}}$, on the jet radius parameter ($R$) for broader area jet cones, going all the way up to $R$ = 1.0. This study presents a comprehensive analysis of high-${p_{T}}$ inclusive jets extending up to 1 TeV to probe the quark-gluon plasma medium at much shorter distance scales. It focuses on quenching effects observed in the quark-gluon plasma formed during Pb-Pb collisions at ${\sqrt{s_{\rm NN}}}$ = 5.02 TeV, particularly for the most-central (0-10%) collisions. Jet-medium interactions represent a pivotal domain of both theoretical and experimental QGP studies, with various models offering different assumptions to describe these phenomena. To illustrate this modular approach, this work computes the nuclear modification factor for inclusive jets via coupling of the MATTER model (which simulates the high virtuality phase of the parton evolution) with the LBT model (which simulates the low virtuality phase of the parton evolution). Additionally, the two successful energy loss models: MARTINI and AdS/CFT are employed to characterize the jet-suppression effectively within the JETSCAPE framework. The results are compared with the experimental data from the ATLAS and CMS detectors, covering jet transverse momentum (${p_{T}}$) ranging from 100 GeV to 1 TeV for ATLAS and 300 GeV to 1 TeV for CMS. The predictions made by the JETSCAPE are consistent in the high ${p_{T}}$ range as well as for extreme jet cone sizes, showing deviation within 10-25%. Our major focus is on calculating the double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=small}}_{\mathrm{AA}}}$) as a function of jet-R and jet-${p_{T}}$, where the experimental results align well with predictions from the JETSCAPE framework, which incorporates the multi-stage evolution of the parton shower and energy loss models.
@@ I Introduction
The extremely hot and dense conditions created at the start of the Big Bang led to what we now understand as the soup of deconfined state of the partons, the quark-gluon plasma (QGP) [1, 2, 3]. The QGP is of such great interest that, to study the properties of this state of matter, experimentalists have spent decades building advanced detectors capable of recreating this extremely dense, soupy state. The Relativistic Heavy-Ion Collider (RHIC) [4, 5] and the Large Hadron Collider (LHC) [6, 7, 8, 9] conduct heavy-ion collisions where the QGP is created for very short periods. During this time, the QGP medium has a substantial impact on the propagation and modification of the parton shower. In particular, the high-${p_{T}}$ jets produced in these heavy ion collisions undergo strong yield suppression and medium modification which are together referred to as jet quenching phenomena [10, 11, 12]. Jets modification has been studied in nucleus-nucleus collisions relative to proton-proton collisions to probe the properties of the QGP via constraints from model-to-data agreement  [13, 14, 15, 16]. The measurement of the nuclear modification factor for jets as well as inclusive jet fragmentation functions has revealed many important characteristics of the quark-gluon plasma [17, 18, 19, 20]. It provides strong confirmation of the interaction of partons with the deconfined plasma, the resulting medium modifications, and the eventual transition to hydrodynamic behavior in the medium [8]. Since there are many conclusive studies based on jet-${R_{AA}}$ [21, 22, 23, 24, 25], our effort here is to push the limits of the current event generators and various energy loss models for a better description of jet quenching phenomena at very high transverse momenta and broader jet cones with the multi-stage evolution of the parton shower using the Jet Energy-loss Tomography with a Statistically and Computationally Advanced Program Envelope (JETSCAPE) framework (version 3.6.5)  [26].
@@
Along with these measurements, we study inclusive jet spectra for p-p and Pb-Pb by varying jet resolution parameters in the anti-${k_{T}}$ algorithm which is realized in the FASTJET software package [27]. The inclusive jet spectrum is of significant interest due to its reduced sensitivity to hadronization effects compared to observables that involve individual final-state hadrons. Here, the area of reconstructed jet cone is defined by the jet radius parameter ${R}$. Thus, by varying ${R}$, the reconstructed jet will include different proportions of energy from the medium response and the quenched jet. We also highlight the discovery of a new sensitivity to the characteristics of the QGP and the fundamental mechanisms of jet quenching in a study for jet yield suppression versus radius ${R}$ [28]. Specifically, the dependence of jet suppression on jet-${R}$, is predicted by theoretical models based on Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence [29] and perturbative QCD [30].
@@
In this paper, we begin with calculating the jet ${R_{AA}}$ for the Pb-Pb collisions in Section II and compare with the experimental data from the ATLAS as well as the CMS detector for a robust test of configuration and the overall JETSCAPE framework. The calculations include the (2+1)D MUSIC [31] model for hydrodynamics, which is ideal for studying various aspects of heavy-ion collisions. Most notably, this work extends to the high-${p_{T}}$ range of jets, up to 1 TeV, which enables us to explore the structure of the QGP medium at much shorter distance scales.
@@
We further exploit the advantages that the JETSCAPE framework offers in Section II.1, that is the coupling of several different energy loss models such as MARTINI [32] and AdS/CFT [33] with the MATTER [34] to explore the quenching effects in a multi-stage manner. This approach provides insight into the interaction and energy loss mechanisms in the medium concerning the virtuality of the partonic jet. Comparing the experimental data with the predictions from these different models, which handle the low virtuality phase, allows us to develop a lucid understanding of the physics governing the above models.
@@
In Section II.2, we proceed with concrete results from the above combinations of several successful models towards the jet radius dependence studies. An important measurement is done by the CMS collaboration with respect to the jet-${R_{AA}}$ in Pb-Pb collisions at ${\sqrt{s_{\rm NN}}}$ = 5.02 TeV at the LHC, for jet cone area covering the radii from 0.2 up to 1  [35]. This measurement gives us an intricate understanding of the behaviour and interaction strength of the high-${p_{T}}$ inclusive collimated jets in proximity to the jet axis, for $R$ = 0.2 and the energy distribution around the cone for values of radii up to 1. The current JETSCAPE framework is based on calculations of perturbative QCD and evolution in a dense medium, thus allowing us to go up to a radius of order 1. We also report the double ratio, i.e., the ratio of ${R_{AA}}$ for a given $R$ with respect to $R$ = 0.2 as a function of jet-${p_{T}}$ and jet radius $R$. This study provides a vivid picture of the energy transactions with the QGP as the jet cone size increases, along with the trends that the JETSCAPE framework predicts.
@@
We conclude this work in Section III, with a concise account of the current JETSCAPE framework’s potential to explain the jet-${R_{AA}}$ for larger area jet cones. We also shed light on the ability of the distinct combination of models to describe the jet-${R_{AA}}$.
@@
I.1 SIMULATION OF ENERGY LOSS USING MULTI-STAGE APPROACH IN JETSCAPE
@@
The JETSCAPE framework provides an ideal environment for carrying out multi-stage energy loss. The hard scattering is generated by Pythia 8 [36] with initial state radiation (ISR) and multiparton interaction (MPI) [37] enabled, and final state radiation (FSR) disabled. For the event wise simulations, the TRENTo model [38] sets up the initial conditions and the viscous hydrodynamic evolution is described by the (2+1)D MUSIC [31] model. This is then followed by Cooper-Frye prescription [39, 40] which converts the fluid cells to hadrons on an isothermal hypersurface [41] at ${T_{\mathrm{SW}}}$= 151 MeV, where ${T_{\mathrm{SW}}}$ is the temperature of the plasma below which particlization [42] occurs at a certain hypersurface. The jet energy loss induced by scattering is calculated in a succession of two stages: MATTER [34, 43] which takes care of the highly virtual phase (the first stage) while the low virtuality phase (the second stage), is handled concurrently by the LBT model [44, 45, 46]. We have also employed the MARTINI [32, 47, 48] and the AdS/CFT model [33] in combination with the MATTER model to explore the low virtuality phase. The virtuality of the parton is defined as ${Q}^{2}=p^{\mu}p_{\mu}-m^{2}$. The parton undergoes energy loss in two stages, when the virtuality of the parton, ${Q^{2}>Q^{2}_{\mathrm{SW}}}$, where ${Q_{\mathrm{SW}}}$ is the switching virtuality, the MATTER model handles the energy loss and the parton is transferred to the LBT model once ${Q^{2}\leq Q^{2}_{\mathrm{SW}}}$. In these calculations, the jet medium interaction includes inelastic medium-induced gluon radiation and a medium recoil. An extensive account of the comparison of the model to the existing jet-${R_{AA}}$ is already done [49], showing that the contribution of medium recoil is quite significant in the modification of jet-${R_{AA}}$ for a complete set of centrality classes ranging from the most central collisions to the peripheral collisions. This version of the JETSCAPE framework [26] encompasses the modifications of a Hard Thermal Loop (HTL) [50] for fixed coupling ($\alpha_{\mathrm{s}}^{\mathrm{fix}}$), running coupling ($\alpha_{\mathrm{s}}$), and with a virtuality dependent factor, ${f(Q^{2})}$, that modulates the effective value of jet transport coefficient ($\hat{q}$). The jet transport coefficient $\hat{q}$ is the ratio of the average of the momentum squared (which is exchanged between the jet parton and the medium) to the unit interaction length traversed by the jet parton in a direction perpendicular to the momentum of the jet parton. The JETSCAPE also accounts for the reduced medium-induced emission in the high virtuality phase, arising from coherence effects. The calculations based on the HTL considering weak-coupling approximation and the limits of high-temperature yield for a $\hat{q}$ is given as [45],
@@
where $C_{a}$ is the representation specific Casimir, $E$ is the energy of the hard parton, and $T$ is the local temperature of the medium. The coherence effects, which reduce the interaction strength of the parton with the medium are also taken into consideration as the virtuality-dependent modulation factor, which regulates the effective value of $\hat{q}$ in the high virtuality MATTER event generator. The virtuality-dependent modulation factor is parameterized as follows [49],
@@
here $Q^{2}$ represents the running virtuality of the hard parton and $\hat{q}^{\mathrm{run}}_{\mathrm{HTL}}$ denotes the formulation with running coupling. Finally, the partons undergo colorless hadronization according to the default Lund string fragmentation from Pythia 8 [36, 51]. The contributions in final jets are from the hard jet shower part and the effect from the soft medium response, the latter is calculated via the Cooper-Frye formula [39, 40]. We reconstruct jets for different radius selections by applying a minimum track-particle transverse momentum requirement of ${p^{\mathrm{track,min}}_{\mathrm{T}}>}$ 0.5 GeV using the anti-${k_{T}}$ algorithm from ${\mathrm{FASTJET}}$, and then compare the results with experimental data. All the parameters involved in the tuning of the constituent models follow the standard set of tunes released by the JETSCAPE Collaboration in an elaborate recent study [49].
Figure 1: (Color online) Differential cross-section of inclusive jets for p+p collisions at ${\sqrt{s_{\rm NN}}}$ = 5.02 TeV with the cone size $R$ = 0.4. Bottom panel shows the ratio of the JETSCAPE to the ATLAS data [24].
Figure 2: Differential cross-section of inclusive jets in Pb+Pb collisions at ${\sqrt{s_{\rm NN}}}$ = 5.02 TeV with the cone size $R$ = 0.4. Bottom panel shows the ratio of the JETSCAPE to the ATLAS data [24].
@@ II Results
This paper covers the collision energy of ${\sqrt{s_{\rm NN}}}$ = 5.02 TeV for the most central (0-10%) Pb-Pb collisions and includes a comparison with selected experimental data from the ATLAS and the CMS collaborations for high-${p_{T}}$ jets. The energy loss depicted in this work is achieved by the coupling of MATTER with LBT model. The secondary LBT model (which handles the low virtuality phase) remains the same throughout until and unless specified. Fig. 1 (top panel) shows the ${p+p}$ collision results for inclusive jet spectra at ${\sqrt{s}=5.02~{}\rm TeV}$ for ${|y_{jet}|<2.8}$, which follow the JETSCAPE PP-19 tune [52], are then compared to the experimental data from the ATLAS [24]. In this study, we have set running coupling constant ($\alpha_{\mathrm{s}}$) to 0.35 for all the calculations.
The ratio of inclusive jet cross-section using the JETSCAPE to the ATLAS data is shown in the bottom panel of Fig. 1, which shows that the results are in the acceptable range. We report the 0-10% most central Pb-Pb jet spectra for $R$ = 0.4 in Fig. 2 (top panel). The ratio of the differential cross-section for inclusive jets in Pb-Pb collisions using JETSCAPE to the ATLAS [24] data are shown in Fig. 2 (bottom panel).
@@
Where ${\sigma_{jet}}$ and ${N_{jet}}$ are the inclusive jet cross-section in ${p+p}$ collisions and the jet yield in Pb+Pb collisions, respectively, which are determined as a function of transverse momentum, ${p_{T}}$ and the jet rapidity, $y_{\mathrm{jet}}$. Moreover, ${N_{evt}}$ represents the number of Pb+Pb collisions within a specific rapidity interval and $\langle T_{\mathrm{AA}}\rangle$ denotes the ratio of the mean number of binary nucleon-nucleon collisions to the inelastic nucleon-nucleon cross-section, which can be determined using the Glauber model of the nuclear collision geometry [53]. The inclusive jet-${R_{AA}}$ is calculated as the ratio of the Pb-Pb and p-p spectra, which is presented in Fig. 3 along with a comparison to the ATLAS data [24]. The JETSCAPE (MATTER + LBT) results are in considerable agreement with the ATLAS measurements. However, a marginal deviation (up to ${18\%}$) is observed in the low ${p_{T}}$ region below 150 GeV, after which the results consistently align with the experimental data as ${p_{T}}$ increases.
These results strongly support the fact that the MUSIC model is not only adequate but also highly effective in describing the experimental observations even at higher jet-${p_{T}}$ in the most central collisions. In contrast to the VISHNU model [54], MUSIC’s hydrodynamic evolution allows for a more accurate representation of the medium’s longitudinal expansion and viscous effects. This leads to a recurring trend of consistent predictions, indicating that MUSIC outperforms other hydrodynamic models in capturing the intricate dynamics of jet interactions within the quark-gluon plasma, especially at high energy scales reaching up to 1 TeV. The MUSIC model shows low sensitivity to coherence effects that dominate at high transverse momentum, compared to VISHNU (2+1)D. This enhances the reliability of MUSIC model for analyzing the jet-suppresions phenomenon at high-${p_{T}}$.
Figure 3: (Color online)the jet-${R_{AA}}$ as a function of jet-${p_{T}}$ for inclusive jets in the most central (0-10%) Pb+Pb collisions at ${\sqrt{s_{\rm NN}}}$ = 5.02 TeV for the jet cone radius $R$ = 0.4 compared with the ATLAS data [24].
Figure 4: The jet-${R_{AA}}$ as a function of jet transverse momentum ($p_{T}^{jet}$) for inclusive jets in the most central (0-10%) Pb+Pb collisions at ${\sqrt{s_{\rm NN}}}$ = 5.02 TeV for the jet cone radii R = 0.2, 0.3, 0.4, 0.6, 0.8 and 1.0 with ${|y_{jet}|<2}$. The plot shows the comparison between the models by red, magenta, and blue markers for (MATTER + LBT), (MATTER + MARTINI), and (MATTER + AdS/CFT), respectively. The predictions are in comparison with CMS data [35] shown with a black marker.